OpenStudy (anonymous):

Please help!! 3n^2-n/n^2-1 divided by n^2/n+1

OpenStudy (anonymous):

\[\frac{ 3n^2 - n }{ n^2 - 1 } \div \frac{ n^2 }{ n +1 }\]

OpenStudy (anonymous):

do u want the ans?

OpenStudy (anonymous):

If you could show me how to work the problem out @physicsboffin that would awesome

OpenStudy (anonymous):

So the answer is 3n-1/n(n-1)?

OpenStudy (anonymous):

can u simplify further?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I have another one would you be able to help?

OpenStudy (anonymous):

yep..quick im gonna go offline

OpenStudy (anonymous):

\[\frac{ 5x +10 }{ x + 2 } * \frac{ 2x }{ 4x - 10 }\]

OpenStudy (compassionate):

Hello, I will be your assistant.

OpenStudy (compassionate):

The first step is to factor your terms and see if anything will cancel. Look at 5x + 10, do you think that will factor? Also, look at 4x - 10.

OpenStudy (anonymous):